how do I help solve for x and y. I only get 0=7??
4x-5y=7
-4x+5y=7

Answer by Alan3354(69443)

4x-5y=7
-4x+5y=7 times -1 --> 4x-5y = -7
-------------
The equations are inconsistent.
4x-5y can't be both 7 & -7.
---
If you graph them, you get parallel lines.
No intersection, no solution.